Tony has two kind of tea, one selling at P280 per kilogram and the other at P305 per kilogram. How many kilograms of each should she get from to make a 25-kilogram blend which will sell at P290 per kilo?
please give me the table for your answers

Answer by nerdybill(7384)

Mixture_Word_Problems/169880 (2008-11-27 06:59:42): Tony has two kind of tea, one selling at P280 per kilogram and the other at P305 per kilogram. How many kilograms of each should she get from to make a 25-kilogram blend which will sell at P290 per kilo?
.
Let x = amount of P280 per kilogram
then
25-x = amount of P305 per kilogram
.
280x + 305(25-x) = 290(25)
280x + 7625 - 305x = 7250
375 - 25x = 0
375 = 25x
15 kilograms = x (amount of P280 per kilogram)
.
amount of P305 per kilogram:
25-x = 25-15 = 10 kilograms